12 If white anhydrous copper(II) sulfate powder is left in the atmosphere it slowly absorbs water vapour giving the blue pent(iv) Determine the value of AH, in kJ mol.
Answer #1

The initial temperature is 19.1 deg C

The final temperature is 26.1 deg C

The temperature change

AT = T, -T; AT = 26.1°C – 19.1°C AT = 7°C

Assume the specific heat of solution to be equal to that of water

C= 4.1801 – 9°C

The mass of solution 3.99 g + 50.0 g = 53.99 g

Irun = WXC XAT Grun = 53.99 g x 4.1801 - 9. 0C x 70 Crun = 1579.8 J

Convert unit from J to kJ

[ 1 ᎬᎫ qrun = 1570.8 Ꭻ x 1000 Ꭻ qrun = 1.57Ꮽ8 kᎫ .

Arxn = -4run qren = -1.5798 kJ

The mass of anhydrous copper(II) sulfate is 3.99 g.

The molar mass of anhydrous copper(II) sulfate is 159.6 g/mol

The number of moles of anhydrous copper(II) sulfate is
3.99 g 159.6 g/mol = 0.025 mol

-1.5798 kJ AH = 0.025 mol AH = -63.2 kJ/mol